A) A cylinder, with a piston pressing down with a constant pressure, is filled with 2.00...

A)

A cylinder, with a piston pressing down with a constant pressure, is filled with 2.00 moles of a gas (n1), and its volume is 41.0 L (V1). If 0.800 mole of gas leak out, and the pressure and temperature remain the same, what is the final volume of the gas inside the cylinder?

Express your answer with the appropriate units

B)

A sample of gas in a cylinder as in the example in Part A has an initial volume of 62.0 L , and you have determined that it contains 1.20 moles of gas. The next day you notice that some of the gas has leaked out. The pressure and temperature remain the same, but the volume has changed to 15.5 L . How many moles of gas (n2) remain in the cylinder?

Express your answer with the appropriate units.
Answer #1

A)

Given:

Vi = 41.0 L

ni = 2.00 mol

nf = 1.20 mol

use:

Vi/ni = Vf/nf

41.0 L / 2.00 mol = Vf / 1.20 mol

Vf = 24.6 L

Answer: 24.6 L

B)

Given:

ni = 1.20 mol

Vi = 62.0 L

Vf = 15.5 L

use:

Vi/ni = Vf/nf

62.0 L / 1.20 mol = 15.5 L / nf

nf = 0.300 mol

Answer: 0.300 mol